OpenStudy (comrad):

a 3.0 kg mass moving to the right at 1.4 m/s collides in a perfectly inelastic collision with a 2.0 kg mass initially at rest. what will the vleocity of the combined mass be after collision? show your work.

OpenStudy (comrad):

Will fan and medal!

OpenStudy (comrad):

@airjordan200

OpenStudy (retireed):

From Wikipedia -- Inelastic collision \[m_a u_a + m_b u_b = \left( m_a + m_b \right) v \, \] Ua is initial velocity of Mass A and Ub is the initial velocity of Mass B where v is the final velocity, which is hence given by \[ v=\frac{m_a u_a + m_b u_b}{m_a + m_b}\] Plug in the values and solve for v...